# BA-LoRA: Bias-Alleviating Low-Rank Adaptation to Mitigate Catastrophic Inheritance in Large Language Models

### v4

@@ 1 Introduction
This phenomenon, termed "Catastrophic Inheritance" by (Chen et al., 2024a), has spurred investigations into mitigation strategies. While constructing less biased datasets and developing more robust model architectures are prominent approaches (Liu & He, 2024), this study explores an alternative: innovations in fine-tuning. Fine-tuning LLMs is a powerful method for enhancing task-specific performance (Han et al., 2024), aligning models with user intent (Ouyang et al., 2022; Xu et al., 2024), and eliciting desired behaviors (Bai et al., 2022; Rafailov et al., 2024). However, fine-tuning large-scale models’ computational and memory demands are substantial (Hu et al., 2021). For instance, 16-bit fine-tuning of a Llama-65B model requires over 780 GB of GPU memory (Dettmers et al., 2024). To address these limitations, parameter-efficient fine-tuning (PEFT) techniques, such as Low-Rank Adaptation (LoRA) (Hu et al., 2021), have gained prominence.
LoRA posits that parameter updates during fine-tuning can be efficiently represented by low-rank matrices. Therefore, given a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, instead of updating all parameters of $W$ directly, LoRA introduces an auxiliary low-rank adapter $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$ with rank $r\ll\min(m,n)$. Here, $A$ and $B$ are learnable matrices initialized as:
  • $A$ is initialized with a scaled normal distribution: $a_{ij}\sim\mathcal{N}(0,\sigma^{2})$.
  • $B$ is initialized to zero: $b_{ij}=0$.
For a given input $X$, the output $Y$ is computed as:
$Y$ $=X(W+\Delta W)=X(W+AB).$ (1)
Only $A$ and $B$ are updated during fine-tuning while $W$ remains frozen. This initialization ensures that $AB=0$ at the start of training, thus preserving the model’s original output. Since the rank $r$ is significantly smaller than the dimensions of $W$, LoRA substantially reduces training overhead compared to full fine-tuning (Hu et al., 2021).
@@
To mitigate the detrimental effects of Catastrophic Inheritance, particularly noise, and imbalance, we propose Bias-Alleviating Low-Rank Adaptation (BA-LoRA). Building upon Principal Singular Values and Singular Vectors Adaptation (PiSSA) (Meng et al., 2024), which addresses convergence issues in standard LoRA, our approach incorporates three distinct regularization terms: a consistency regularizer, a diversity regularizer, and a singular value decomposition (SVD) regularizer. The consistency regularizer preserves valuable pre-trained knowledge during fine-tuning, while the diversity regularizer encourages varied model outputs. The SVD regularizer enhances the generalization capabilities of generative models. Recognizing the fundamental differences between Natural Language Understanding (NLU) and Natural Language Generation (NLG), such as determinism in NLU versus diversity in NLG, we tailor our regularization strategies accordingly.
To evaluate the efficacy of BA-LoRA, we conduct comprehensive experiments across diverse benchmarks, including mathematical reasoning (GSM8K (Cobbe et al., 2021) and MATH (Yu et al., 2023)), coding (HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021)), natural language understanding (GLUE (Wang et al., 2018)), and general language evaluation (MT-Bench (Zheng et al., 2024)). Our experiments utilize prominent LLMs such as LLaMA 2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), and Gemma-7B (Team et al., 2024), as well as encoder-only architectures like RoBERTa-large (Liu et al., 2019) and DeBERTa-v3-base (He et al., 2021b). The results unequivocally demonstrate BA-LoRA’s superiority over LoRA and its state-of-the-art variants. Moreover, the extended experiments demonstrate that our method effectively attenuates noise inherited from pre-training, leading to more robust and generalizable models.
@@ 3.2.3 Regularizations for NLG Tasks
where $\mathcal{P}_{\mathrm{pt}}(y_{t}\mid y_{<t},x)$ and $\mathcal{P}_{\mathrm{ft}}(y_{t}\mid y_{<t},x)$ represent the conditional probability distributions of the pre-trained model and the fine-tuned model, respectively. For the current token $y_{t}$, given the input $x$ and the preceding token sequence $y_{<t}$. KLD encourages the model to continuously retain useful pre-training information during the fine-tuning process, which is crucial for maintaining the style and coherence of the generation task.
@@
Singular Value Decomposition Regularization To enhance the generalization of generative models while ensuring computational efficiency, we introduce a regularization technique based on randomized SVD (Pasand & Bashivan, ). This method efficiently approximates the dominant singular values of a matrix, significantly reducing the computational and memory costs associated with traditional SVD while preserving key information about the data.
@@ 4.2 Implementation Details
In our experiments, we adopt the PiSSA (Meng et al., 2024) implementation strategy. We compute the loss using only the responses from the instruction-following dataset, ensuring lora_dropout to 0. We utilize the Float32 computation type for both the base model and the adapter in BA-LoRA. For the NLU tasks, we set the hyperparameters as: $k=5$, $\lambda_{1}=1e-1$, $\lambda_{2}=4e-2$, and $\lambda_{3}=1e-2$. We set lora_r $=$ lora_alpha $=128$ and use AdamW (Loshchilov & Hutter, 2017) optimizer with a batch size of 128, a learning rate of $52e-$ 5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. For the NLG tasks, the hyperparameters are set as: $k=10$, $\lambda_{1}=1e-4$, $\lambda_{2}=3e-1$, and $\lambda_{3}=1e-1$. We set lora_r as 8 and select lora_alpha in 8, 16. We utilize AdamW with a linear learning rate schedule to optimize and tune the learning rate (LR) from $41e-$ 4, $42e-$ 4, $43e-$ 4, $44e-$ 4, $45e-$ 4, $46e-$ 4, $55e-$ 5, $53e-$ 5. Batch sizes (BS) are selected from 6, 8, 16, 32. Appendix Section B presents the detailed hyperparameters we utilized on the GLUE benchmark. comparison. All experiments were conducted using NVIDIA A40 (48G) GPUs. The results were derived from three independent experiments conducted under identical conditions, which were averaged to ensure consistency, robustness, and variability mitigation.
@@ 4.3.1 Analysis of the NLG and NLU Performance of BA-LoRA
To evaluate BA-LoRA’s effectiveness on NLG tasks, we fine-tuned LLaMA-2-7B, Mistral-7B, and Gemma-7B on the MetaMathQA (Yu et al., 2023) and assessed their mathematical problem-solving capabilities using the GSM8K (Cobbe et al., 2021) and MATH (Yu et al., 2023) validation sets, reporting Accuracy. Similarly, models were fine-tuned on the CodeFeedback (Zheng et al., 2024) and evaluated for coding proficiency via HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), with PASS@1 metrics reported. To assess conversational abilities, models were trained on the WizardLM-Evol-Instruct (Xu et al., 2024) and evaluated on MT-Bench (Zheng et al., 2024), with response quality judged by GPT-4 and first turn scores reported. All experiments utilized 100K data points and a single training epoch for efficiency.
Table 1 presents the experimental outcomes, clearly demonstrating BA-LoRA’s superior performance compared to baseline methods. For instance, BA-LoRA enhanced LLaMA 2-7B, Mistral-7B, and Gemma-7B performance on GSM8K by 1.82%, 1.14%, and 0.55%, respectively, compared to PiSSA. HumanEval improvements were 2.03%, 1.11%, and 1.26%, while MT-Bench enhancements reached 0.24%, 0.18%, and 0.07%. Notably, BA-LoRA achieved a remarkable 6.92% performance uplift over full parameter fine-tuning on Gemma, utilizing only 2.3% of trainable parameters across five tasks.
To assess the effectiveness of BA-LoRA on natural language understanding (NLU) tasks, we conducted experiments on the GLUE benchmark (Wang et al., 2018), which includes two single-sentence classification tasks (CoLA, SST), five paired-text classification tasks (MNLI, RTE, QQP, MRPC, QNLI), and one text similarity prediction task (STS-B). The evaluation metrics comprise the overall matched and mismatched accuracy for MNLI, the Matthews correlation coefficient for CoLA, the Pearson correlation coefficient for STS-B, and accuracy for the remaining tasks. We used the DeBERTa-v3-base model (He et al., 2021b) and compared BA-LoRA against ten baseline methods, including Full Fine-Tuning (Full FT), BitFit (Zaken et al., 2021), HAdapter (Houlsby et al., 2019a), PAdapter (Pfeiffer et al., 2020), LoRA (Hu et al., 2021), LoHA (Hyeon-Woo et al., 2021), DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2023a), and PiSSA (Meng et al., 2024).
@@
A comparative analysis of Tables 1 and 2 reveals BA-LoRA’s consistent performance advantages across both NLG and NLU tasks. This indicates BA-LoRA’s proficiency in augmenting both generative and comprehension capabilities for language models. By incorporating consistency, diversity, and SVD regularization, BA-LoRA effectively mitigates the adverse effects of Catastrophic Inheritance, fostering consistent, diverse, and generalized model outputs. Furthermore, BA-LoRA’s modest computational requirements render it suitable for efficient fine-tuning of LLMs with limited resources.
4.4 Evaluating the Performance of Different Ranks
We compare the performance of BA-LoRA, LoRA, and PiSSA at different ranks using LLaMA-2-7B and Mistral-7B-v0.1 models. Each method is fine-tuned for one epoch on the MetaMathQA-100K dataset with ranks ranging from 1 to 128 and evaluated on the GSM8K and MATH. As Figure  1 shows, BA-LoRA consistently outperforms LoRA and PiSSA across all rank settings and datasets. As the rank increases, the performance of BA-LoRA and PiSSA surpasses full parameter fine-tuning. However, BA-LoRA performs better, especially on Mistral-7B-v0.1.
@@ 4.4.2 Analysis on Mitigating Imbalanced Data
This experiment evaluates the effectiveness of BA-LoRA in addressing imbalanced data. Specifically, using the MNLI dataset, LoRA and BA-LoRA are applied to fine-tune the BERT-L and GPT-2-XL models, respectively. The hidden layer features of the last training step are extracted and visualized using t-SNE (Van der Maaten & Hinton, 2008) technology for comparison.
@@ 5 Conclusion
This paper introduces BA-LoRA, a novel parameter-efficient fine-tuning method designed to mitigate catastrophic inheritance in pre-trained language models. BA-LoRA incorporates three key components: consistency regularization, diversity regularization, and singular value decomposition regularization. These components work in concert to preserve pre-training knowledge, enhance output diversity, and improve model generalization. Extensive experiments demonstrate that BA-LoRA consistently outperforms existing baselines on various NLG and NLU tasks while robust to noisy and imbalanced pre-training data. Furthermore, our ablation studies confirm the effectiveness of the three regularization terms both individually and in combination. These results highlight BA-LoRA’s potential as a general-purpose fine-tuning method for pre-trained language models and effectively address the key challenges of deploying these models in real applications.
@@ A.3 Examples of noise in pre-training data
Duplicity: The presence of identical or similar content in the data can lead to overfitting and privacy leakage risks (Elazar et al., 2023), (Carlini et al., 2022), (Hernandez et al., 2022). Corruption/noise: Inconsistent or erroneous inputs in the training data can affect model robustness and downstream task performance (Elazar et al., 2023), (Fan et al., 2024), (Caswell et al., 2021). Contamination: Leakage of the test set into the training set may lead to distorted model evaluation results (Roberts et al., 2023), (Schaeffer, 2023), (Jiang et al., 2024b).
@@
Category imbalance: Too few samples of certain categories cause the model to perform poorly in predicting these categories, producing bias (Xu et al., 2023a), (Zhu et al., 2024), (Parashar et al., 2024).
@@
Toxic and harmful content: The training data may contain offensive, biased, or harmful content that may cause the model to generate harmful or inappropriate outputs (Zou et al., 2023), (Sun et al., 2024).
@@ C.2 Impact of Regularization on LoRA Variants Performance
C.3 t-SNE Visualizations of Feature Evolution during the Fine-tuning with LoRA and BA-LoRA
@@ D.1 Computational Cost Analysis
To address these challenges, we introduce a randomized SVD approach, which achieves significant computational savings while preserving sufficient accuracy. Randomized SVD leverages random projections to approximate the dominant singular values and vectors, enabling faster computation and enhanced scalability for large matrices. This optimization ensures that computational demands remain manageable, even for models with high-dimensional weight matrices, thereby supporting scalability across diverse architectures.
To assess the practical implications of our approach, we conducted experiments measuring the average training time and memory usage across multiple runs. Specifically, we fine-tuned the DeBERTa-v3-base model on the full GLUE benchmark using an NVIDIA A40 GPU. The results for three PEFT methods—LoRA, PiSSA, and our proposed BA-LoRA—are summarized below:
  • LoRA: Averaged 14.45 hours of training time with an average memory usage of 6.3 GB.
  • PiSSA: Averaged 12.07 hours of training time with an average memory usage of 6.5 GB.
  • BA-LoRA: Averaged 15.82 hours of training time with an average memory usage of 7.1 GB.
Although BA-LoRA incurs marginally higher computational costs compared to LoRA and PiSSA, these increases stem primarily from its advanced regularization mechanisms, including the innovative use of randomized SVD. Crucially, these mechanisms effectively mitigate the catastrophic forgetting problem, resulting in substantial performance improvements. Considering the trade-off between computational overhead and model performance, the additional costs of BA-LoRA are both manageable and well-justified, especially in scenarios where achieving superior accuracy is paramount.

### v5

@@ 1 Introduction
This phenomenon, termed “Catastrophic Inheritance” by (Chen et al., 2024a), has spurred investigations into mitigation strategies. While constructing less biased datasets and developing more robust model architectures are prominent approaches (Liu & He, 2024), this study explores an alternative: innovations in fine-tuning. Fine-tuning LLMs is a powerful method for enhancing task-specific performance (Han et al., 2024), aligning models with user intent (Ouyang et al., 2022; Xu et al., 2024), and eliciting desired behaviors (Bai et al., 2022; Rafailov et al., 2024). However, fine-tuning large-scale models’ computational and memory demands are substantial (Hu et al., 2021). For instance, 16-bit fine-tuning of a Llama-65B model requires over 780 GB of GPU memory (Dettmers et al., 2024). To address these limitations, parameter-efficient fine-tuning (PEFT) techniques, such as Low-Rank Adaptation (LoRA) (Hu et al., 2021), have gained prominence.
LoRA enables efficient fine-tuning of large pre-trained models by approximating parameter updates using low-rank matrices. As illustrated in Figure 1a, Full Fine-tuning involves directly updating the entire weight matrix $W$. In contrast, LoRA (shown in Figure 1b) introduces a learnable low-rank adapter $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$ are trainable low-rank matrices with a user-specified rank $r\ll\min(m,n)$. This ensures that only $A$ and $B$ are updated during training, while the original weights $W$ remain frozen. By initializing $A$ with scaled random values and $B$ to zero, LoRA preserves the pre-trained model’s initial output at the start of training. For an input $X$, the output is computed as $Y=X(W+AB)$, ensuring consistency with the original model. This method significantly reduces computational costs by limiting the number of trainable parameters, making it a practical solution for adapting large models (Hu et al., 2021).
@@
To mitigate the detrimental effects of Catastrophic Inheritance, particularly noise, and imbalance, we propose Bias-Alleviating Low-Rank Adaptation (BA-LoRA). As depicted in Figure 1c, building upon Principal Singular Values and Singular Vectors Adaptation (PiSSA) (Meng et al., 2024), which addresses convergence issues in standard LoRA, BA-LoRA incorporates three distinct regularization terms: a consistency regularizer, a diversity regularizer, and a singular value decomposition (SVD) regularizer. The consistency regularizer preserves valuable pre-trained knowledge during fine-tuning, while the diversity regularizer encourages varied model outputs, and the SVD regularizer enhances the generalization capabilities of generative models. Recognizing the fundamental differences between Natural Language Understanding (NLU) and Natural Language Generation (NLG), such as determinism in NLU versus diversity in NLG, we tailor our regularization strategies accordingly.
Figure 1: Framework comparisons of Full Fine-tuning (a), LoRA (b), and BA-LoRA (c). Where the blue modules indicate frozen parameters during training, the orange modules denote trainable components, and the dark red module indicates the custom regularization terms.
@@ 3.2.3 Regularizations for NLG Tasks
where $\mathcal{P}_{\mathrm{pt}}(y_{t}\mid y_{<t},x)$ and $\mathcal{P}_{\mathrm{ft}}(y_{t}\mid y_{<t},x)$ represent the conditional probability distributions of the pre-trained model and the fine-tuned model, respectively. For the current token $y_{t}$, given the input $x$ and the preceding token sequence $y_{<t}$. KLD regularization encourages the model to continuously retain useful pre-training information during the fine-tuning process, which balances the trade-off between maintaining pre-trained knowledge and optimizing downstream task performance.
@@
Singular Value Decomposition Regularization
To enhance the generalization of generative models while ensuring computational efficiency, we introduce a regularization technique based on randomized SVD (Pasand & Bashivan, ). This method efficiently approximates the dominant singular values of a matrix, significantly reducing the computational and memory costs associated with traditional SVD while preserving key information about the data.
@@ 4.2 Implementation Details
In our experiments, we adopt the PiSSA (Meng et al., 2024) implementation strategy. We compute the loss using only the responses from the instruction-following dataset, ensuring lora_dropout to 0. We utilize the Float32 computation type for the base model and the adapter in BA-LoRA. LoRA adapters are applied to all linear layers. For the NLU tasks, we set the hyperparameters as: $k=5$, $\lambda_{1}=1e-1$, $\lambda_{2}=1e-2$, and $\lambda_{3}=1e-2$. We set lora_r $=$ lora_alpha $=128$ and use AdamW (Loshchilov & Hutter, 2017) optimizer with a batch size of 128, a learning rate of $52e-$ 5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. For the NLG tasks, the hyperparameters are set as: $k=10$, $\lambda_{1}=1e-5$, $\lambda_{2}=1e-1$, and $\lambda_{3}=1e-1$. We set lora_r as 8 and select lora_alpha in 8, 16. We utilize AdamW with a linear learning rate schedule to optimize and tune the learning rate (LR) from $41e-$ 4, $42e-$ 4, $43e-$ 4, $44e-$ 4, $45e-$ 4, $46e-$ 4, $55e-$ 5, $53e-$ 5. Batch sizes (BS) are selected from 6, 8, 16, and 32. Appendix Section B presents the detailed hyperparameters we utilized on the GLUE benchmark. comparison. All experiments were conducted using NVIDIA A40 (48G) GPUs. The results were derived from three independent experiments conducted under identical conditions, which were averaged to ensure consistency, robustness, and variability mitigation.
@@ 4.3.1 Analysis of the NLG and NLU Performance of BA-LoRA
To evaluate BA-LoRA’s effectiveness on NLG tasks, we fine-tuned LLaMA-2-7B, Mistral-7B, and Gemma-7B on the MetaMathQA (Yu et al., 2023) and assessed their mathematical problem-solving capabilities using the GSM8K (Cobbe et al., 2021) and MATH (Yu et al., 2023) validation sets, reporting Accuracy. Similarly, models were fine-tuned on the CodeFeedback (Zheng et al., 2024) and evaluated for coding proficiency via HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), with PASS@1 metrics reported. To assess conversational abilities, models were trained on the WizardLM-Evol-Instruct (Xu et al., 2024) and evaluated on MT-Bench (Zheng et al., 2024), with response quality judged by GPT-4 and first-turn scores reported. All experiments utilized 100K data points and a single training epoch for efficiency.
The experimental results in Table 1 demonstrate the superior performance of BA-LoRA over the baseline methods. For example, BA-LoRA enhanced LLaMA 2-7B, Mistral-7B, and Gemma-7B performance on GSM8K by 1.82%, 1.14%, and 0.55%, respectively, compared to PiSSA. HumanEval improvements were 2.03%, 1.11%, and 1.26%, while MT-Bench enhancements reached 0.24%, 0.18%, and 0.07%. Notably, BA-LoRA achieved a remarkable 6.92% performance uplift over full parameter fine-tuning on Gemma, utilizing only 2.3% of trainable parameters.
To assess the effectiveness of BA-LoRA on natural language understanding (NLU) tasks, we conducted experiments on the GLUE benchmark (Wang et al., 2018), which includes two single-sentence classification tasks (CoLA, SST), five paired-text classification tasks (MNLI, RTE, QQP, MRPC, QNLI), and one text similarity prediction task (STS-B). The evaluation metrics comprise the overall matched and mismatched accuracy for MNLI, the Matthews correlation coefficient for CoLA, the Pearson correlation coefficient for STS-B, and the accuracy for the remaining tasks. We used the DeBERTa-v3-base model (He et al., 2021b) and compared BA-LoRA against ten baseline methods, including Full Fine-Tuning (Full FT), BitFit (Zaken et al., 2021), HAdapter (Houlsby et al., 2019a), PAdapter (Pfeiffer et al., 2020), LoRA (Hu et al., 2021), LoHA (Hyeon-Woo et al., 2021), DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2023a), and PiSSA (Meng et al., 2024).
@@
A comparative analysis of Tables 1 and 2 reveals BA-LoRA’s consistent performance advantages across NLG and NLU tasks. This indicates BA-LoRA’s proficiency in augmenting both generative and comprehension capabilities for language models. By incorporating consistency, diversity, and SVD regularization, BA-LoRA effectively mitigates the adverse effects of Catastrophic Inheritance, fostering consistent, diverse, and generalized model outputs. Furthermore, BA-LoRA’s modest computational requirements render it suitable for efficient fine-tuning of LLMs with limited resources.
@@ 4.3.3 Analysis on Mitigating Imbalanced Data
This experiment evaluates the effectiveness of BA-LoRA in addressing imbalanced data. Specifically, using the MNLI task of the GLUE benchmark, LoRA, and BA-LoRA are applied to fine-tune the BERT-L and GPT-2-XL models, respectively. The hidden layer features of the last training step are extracted and visualized using t-SNE (Van der Maaten & Hinton, 2008) technology for comparison.
@@ 5 Conclusion
This paper introduces BA-LoRA, a novel parameter-efficient fine-tuning method designed to mitigate catastrophic inheritance in pre-trained language models. BA-LoRA incorporates three key components: consistency regularization, diversity regularization, and singular value decomposition regularization. These components work in concert to preserve pre-training knowledge, enhance output diversity, and improve model generalization. Extensive experiments demonstrate that BA-LoRA consistently outperforms existing baselines on various NLG and NLU tasks while being robust to noisy and imbalanced pre-training data. Furthermore, our ablation studies confirm the effectiveness of the three regularization terms. These results highlight BA-LoRA’s potential as a general-purpose fine-tuning method for pre-trained language models and effectively address the key challenges of deploying these models in real applications.
@@ A.3 Examples of noise in pre-training data
Duplicity: The presence of identical or similar content in the data can lead to overfitting and privacy leakage risks (Elazar et al., 2023; Carlini et al., 2022; Hernandez et al., 2022).
Corruption/noise: Inconsistent or erroneous inputs in the training data can affect model robustness and downstream task performance (Elazar et al., 2023; Fan et al., 2024; Caswell et al., 2021).
Contamination: Leakage of the test set into the training set may lead to distorted model evaluation results (Roberts et al., 2023; Schaeffer, 2023; Jiang et al., 2024b).
@@
Category imbalance: Too few samples of certain categories cause the model to perform poorly in predicting these categories, producing bias (Xu et al., 2023a; Zhu et al., 2024; Parashar et al., 2024), (Zhu et al., 2024).
@@
Toxic and harmful content: The training data may contain offensive, biased, or harmful content that may cause the model to generate harmful or inappropriate outputs (Zou et al., 2023; Sun et al., 2024).
@@ C.2 Impact of Regularization on LoRA Variants Performance
C.3 Evaluating the Performance of Different Ranks
We compare the performance of BA-LoRA, LoRA, and PiSSA at different ranks using LLaMA-2-7B and Mistral-7B-v0.1 models. Each method is fine-tuned for one epoch on the MetaMathQA-100K dataset with ranks ranging from 1 to 128 and evaluated on the GSM8K and MATH. As Figure  5 shows, BA-LoRA consistently outperforms LoRA and PiSSA across all rank settings and datasets. As the rank increases, the performance of BA-LoRA and PiSSA surpasses full parameter fine-tuning. However, BA-LoRA performs better, especially on Mistral-7B-v0.1.
C.4 Hyperparameter Analysis
To evaluate the stability of hyperparameters through sensitivity analysis, we systematically tested $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ on LLaMA-2-7B (Touvron et al., 2023) using the GSM8K (Cobbe et al., 2021) dataset. By varying $\lambda_{1}$ in $[0.0,0.1]$ and $\lambda_{2}/\lambda_{3}$ in $[0.01,0.2]$ (with other parameters fixed).
The results (Figure 6) show that $\lambda_{1}=0.00001$ minimizes interference with the pre-trained model, while $\lambda_{2}=0.1$ and $\lambda_{3}=0.1$ maximize regularization effectiveness. Increasing $\lambda_{1}$ to $0.001$ or $0.01$ reduces performance, and variations in $\lambda_{2}$ and $\lambda_{3}$ (e.g., $0.05$ or $0.15$) lead to suboptimal performance. This confirms that the selected hyperparameters provide the best trade-off for model performance.
Figure 6: Sensitivity Analysis of Regularization Coefficients $(\lambda_{1},\lambda_{2},\lambda_{3})$ in $\mathcal{L}_{\mathrm{task\_NLG}}$ on LLaMA-2-7B Accuracy for GSM8K.
C.5 t-SNE Visualizations of Feature Evolution during the Fine-tuning with LoRA and BA-LoRA
@@ D.1 Computational Cost Analysis
Table 10: Computational cost analysis of different fine-tuning methods.
Method #Params Memory Cost Training Time GSM8K MATH
Full FT 6738M >48 GB >24h $49.13_{\pm 0.21}$ $7.29_{\pm 0.22}$
LoRA 320M 37.81 GB 5h 32min $42.47_{\pm 0.29}$ $5.60_{\pm 0.35}$
PiSSA 320M 33.94 GB 4h 28min $52.37_{\pm 0.52}$ $7.76_{\pm 0.19}$
BA-LoRA 320M 40.68 GB 5h 41min $54.53_{\pm 0.41}$ $9.21_{\pm 0.17}$
@@
To reconcile this tension between computational efficiency and methodological rigor, our design incorporates targeted optimizations at both the algorithmic and implementation levels. First, during the design phase, we specifically considered and addressed this issue and implemented strategies to minimize resource consumption without sacrificing performance. As described in Section 3.2.3 "Singular Value Decomposition Regularization" of the main text: To enhance the generalization of generative models while ensuring computational efficiency, we introduce a regularization technique based on randomized SVD (Pasand & Bashivan, ). This method efficiently approximates the dominant singular values of a matrix, significantly reducing the computational and memory costs associated with traditional SVD while preserving key information about the data.
Second, to further demonstrate the effectiveness of our method and its limited additional resource consumption, we calculated the average training time and average memory usage across the first three training runs. Specifically, we selected the NLG task (which involves high training resource consumption) for experimental testing. We compared the training parameters, memory usage, and training time across four methods: Full FT, LoRA, PiSSA, and BA-LoRA. Additionally, we evaluated the fine-tuning performance on the GSM8K and MATH benchmarks. The experiments were conducted on an A40 GPU using DeepSpeed (Rasley et al., 2020) ZeRO-2 stage optimization. The LLaMA-2-7B model was fine-tuned on the first 100,000 entries of the MetaMathQA dataset.
As shown in Table 10, BA-LoRA maintains the same number of fine-tuned parameters (320M) as standard LoRA and PiSSA when applied to the LLaMA-2-7B model for NLG tasks. Furthermore, while BA-LoRA incurs marginally higher computational costs than LoRA and PiSSA, this increase primarily stems from its advanced regularization mechanisms (e.g., randomized SVD). Crucially, these mechanisms effectively mitigate catastrophic forgetting while delivering substantial performance improvements. The additional computational overhead remains manageable and well-justified, particularly in scenarios where superior accuracy is paramount. The trade-off between computational cost and performance demonstrates BA-LoRA’s practical viability.